answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

When 2 dice are rolled what is the probability of getting a sum greater than 9?

Each die can come up in 6 different ways.
There are (6 x 6) = 36 possible results.

The results that show greater than 9 are:

4 and 6, 6 and 4

5 and 6, 6 and 5

6 and 6, 6 and 6

Six possibilities.

The probability of a sum greater than 9 is [ 6 / 36 ] = 1/6th = 16 and 2/3 percent.

(Odds are 5 to 1 against.)

Features on the probability scale?

well the question you wrote do not make any and please stop wasting your time on question like these be more specific you retard!

Find the probability of rolling a sum of 10 with 2 dice?

Well, there are 36 options to be rolled. The only combos that will add up to ten are 5+5 and 6+4. So that's 2 options out of 36. Hence 2/36.

simplified to 1/18 OR 5.56%

What is the probability of obtaining a number that is odd and smaller than 5?

If you are selecting from any number than virtually 0. If you are rolling a die then there are 6 possible outcomes, and 2 numbers that are odd and less than 5 so 2/6=1/3.

What is the theoretical probability of tossing two fair coins and having one land on heads an the other one lands on tails?

Depending on the specifications, it could be 1/4 (25%) or 1/2 (50%). If you specify that Coin A must be heads and Coin B is tails, then it is 25%: Because you have 1/2 chance for each coin, you multiply these chances together : (1/2) x (1/2) = 1/4

However if you just want 'one of the coins' to be heads, and 'the other coin' to be tails, then it is 1/2 (50%). Here are the 4 possible outcomes:

  1. Both A & B are heads
  2. Both A & B are tails
  3. A is heads, B is tails
  4. A is tails, B is heads.

So for outcomes # 3 & 4 meets our criteria. This is 2 out of the 4 possible, or 2/4 = 1/2 or 50%.

What is the probability that a coin lands on its heads and tails?

The probability of the coin flip being heads or tails is 100%.

What percent of genetic makeup is shared by identical twins?

100%

Identical twins develop from the same egg, so have the same DNA. The only thing different between identical twins is finger prints, and retinal patterns.

What does accountability mean?

Accountability is synonymous with responsibility.
Means you answer to someone in regards to job performance/ your actions.

What is the definition of a probabilistic system?

A probabilistic system is one that is understandable, indirectly, using probability analysis.

Related Information:

In these systems, specific outcomes can't be predicted with precision, but the probabilities of given outcomes are known.

What is the probability to roll a 7 or 8 on a eight-sided die?

It depends on what numbers are on its faces.

If it is a fair die with the numbers 1 to 8 - and there is no reason to assume so - the answer is 1/4 or 0.25

What is the probability of the birthday of 6 students fall in exactly 2 calendar months is?

It is not possible to answer the question since births are not uniformly distributed over the year. Furthermore, calendar months are not the same length, either.

What is the probability of getting a king?

What part of the world are you in? If you're talking about the united states replacing the president with a king, that is impossible. There is no probability of the united states getting a king.

If there is a bag that contains marbles numbered 1-10 what is the probability that you would get a 9 then a 3?

If you don't put the first one back in the bag before the second draw, then the total
possible number of draws is (10 x 9) = 90, and only one of them is successful. So the
probability is 1/90 = 0.0111 = 1.11% (rounded).

If you do put the first one back in the bag before the second draw, then the total possible
number of draws is (10 x 10) = 100, and only one of them is the one you want. So the
probability is 1/100 = 0.01 = 1%.

You and your boyfriend are having a baby and he is a twin and his twin brother is having twin girls are you going to have twins?

Twins normally skip a generation... but there is no telling until you get a sonogram.

if you DO only have one child at this time, that child has the twin gene, and could have twins.

its the women who determs if you are having twins

Trending Questions
What is the probability of drawing 2 cards that are all kings from a deck of cards? How many numbers used for 10000 permutation? How do you get the negative of 4 to the 7th power? What are the Eye defects and their correction? If the probability of hitting a target when throwing a dart is five out of seven and if six darts are thrown what is the probability of getting exactly four hits? What is the probability you will roll two prime numbers on a dice? Which event occurs in telophase II? Importances of probability in real life? What was the name of the area used by the Indians? What is the probability of picking a month that starts with J given that it ends in y? In ace combat 5 can you save chopper? How do you do probability saying no? If you roll a 6 faced dice how may outcomes are there? What is the probability of getting five questions correct on a 20 question multiple choice test? How many times do you have to shuffle a deck of cards so that you will win solitare? If On a Saturday evening 34 of the people in Chicago go out to dinner 18 see a movie 13 have a party and 35 stay home If seven people are randomly selected what is the probability that one eat? Four different sets of objects contain 2 4 6 and 7 objects respectively. How many unique combinations can be formed by picking one object from each set? What is the probability of choosing something to wear? What is the probability of rolling a three or a six on a die then tossing a heads on a coin? How do you do probability with dice?